# Random Utility Models with Skewed Random Components:

### v1

@@ 3.4 Patterns of Substitution and the Lack of IIA
We note in passing that IIA also does not hold for standard multinomial probit models, i.e., the multinomial probit with iid normal random utility components. See Section S.C in the supplementary appendix for details.
@@ 4.2 On the Estimation of Parameter Ratios
Assume that (i) $\hat{\beta}_{\mathrm{Q}}\rightarrow^{p}\beta_{\mathrm{Q}}^{\ast}\in\mathcal{B}$, and $\beta_{\mathrm{Q}}^{\ast}$ is the unique solution to $E\left[X_{i}^{\prime}\Lambda_{i}^{\mathrm{Q}}\left(X_{i}\beta\right)\overrightarrow{Y}_{i}\right]=0;$
@@ 8 Conclusions and Further Research Directions
Third, there are various other flavors of logit models used by economists. The nested logit explicitly allows a correlation between different subsets of alternatives that are uncorrelated within a nest. Exploring the properties of the SEVI analogue would be of interest, given the quite different pattern of substitution within a nest. The generalized multinomial logit’s gamma parameter determines how the random component of the mixed logit interacts with the scale parameter (Fiebig et al., (2010)). It is unclear how this tradeoff is influenced by a SEVI or LEVI core. Another interesting question is how the cut-off points and parameter estimates are influenced in an ordered logit model when switching its foundation from LEVI to SEVI error components. As discussed in Section 3.3, the SEVI analogue of the LEVI-type model for completely rank-ordered can be developed, but the exact details warrant further studies.
@@ 9 Appendix of Proofs
Proof of Theorem 1. Define the event
$\mathcal{E}_{k,j}:v_{ij}+\varepsilon_{ij}\geq v_{ik}+\varepsilon_{ik}$
for each $k\in\mathcal{J}_{-j}.$ Equivalently,
$\mathcal{E}_{k,j}:-v_{ij}+\tilde{\varepsilon}_{ij}\leq-v_{ik}+\tilde{\varepsilon}_{ik}.$
@@
$=1-\sum_{k_{1}:\left\{k_{1}\right\}\subseteq\mathcal{J}_{-j}}\Pr\left(\mathcal{E}_{k,j}^{c}|v_{i}\right)+\sum_{k_{1}<k_{2},\left\{k_{1},k_{2}\right\}\subseteq\mathcal{J}_{-j}}\Pr\left(\mathcal{E}_{k_{1},j}^{c}\cap\mathcal{E}_{k_{2},j}^{c}|v_{i}\right)$
$+\ldots+\left(-1\right)^{J-1}\sum_{k_{1}<\ldots<k_{J-1},\left\{k_{1},\ldots,k_{J-1}\right\}\subseteq\mathcal{J}_{-j}}\Pr\left(\mathcal{E}_{k_{1,j}}^{c}\cap\mathcal{E}_{k_{2},j}^{c}\cap\ldots\cap\mathcal{E}_{k_{J-1},j}^{c}|v_{i}\right),$ (21)
where we have used the inclusion-exclusion principle (e.g., Proposition 4.4 of Ross, (2013)) for computing $\Pr(\cup_{{}_{k\in\mathcal{J}_{-j}}}\mathcal{E}_{k,j}^{c}|v_{i})$, the probability of a union of $\left(J-1\right)$ events. Note that $\mathcal{E}_{k_{1},j}^{c}$ and $\mathcal{E}_{k_{2},j}^{c}$ are not independent.
@@
$-v_{ij}+\tilde{\varepsilon}_{ij}\geq-v_{ik}+\tilde{\varepsilon}_{ik}\text{ for all }k\in\left\{k_{1},\ldots,k_{\ell}\right\}\subseteq\mathcal{J}_{-j}.$
The probability of the above event is the same as the probability of choosing the $j$-th alternative out of the set of alternatives $\left\{k_{1},\ldots,k_{\ell},j\right\}$ in the standard LEVI model with $\left\{-v_{i,k_{1}},\ldots,-v_{i,k_{\ell}},-v_{i,j}\right\}$ as the systematic utilities. Hence, this probability is equal to
@@
Proof of Proposition 2. The first result is the same as that given in Theorem 1.
To prove the second result, we note that
@@
Proof of Proposition 3. Recall that $G\left(\cdot\right)$ and $g\left(\cdot\right)$ are the CDF and pdf of the LEVI distribution. For $\varepsilon_{j}\thicksim iid$ $\mathrm{SEVI}$ and $\tilde{\varepsilon}_{j}\thicksim iid$ $\mathrm{LEVI},$ we have
@@
$=\sum_{j=1}^{J}\left[v_{j}-E\left(\tilde{\varepsilon}_{j}|v_{k}-v_{j}+\tilde{\varepsilon}_{j}\leq\tilde{\varepsilon}_{k}\text{, }k\in\mathcal{J}_{-j}\right)\right]P^{\mathrm{SEVI}}\left(Y=j|v\right)$
@@
By employing a change of variable and letting
@@
$=\left[\gamma+v_{j}+\ln\left(\sum_{k\in\mathcal{S}_{\ell}\cup j}\exp\left(-v_{k})\right)\right)\right]\frac{\exp\left(-v_{j}\right)}{\sum_{k\in\mathcal{S}_{\ell}\cup j}\exp\left(-v_{k})\right)}.$
where the second-to-last line follows because for any $\xi>0,$ we can show that
@@
For notational simplicity, let $\mathcal{S}_{\ell,j}:=\mathcal{S}_{\ell}\cup j.$ Then
@@
$=-\sum_{j=1}^{J}\sum_{\ell=1}^{J-1}\left(-1\right)^{\ell}\sum_{k_{1}<\ldots<k_{\ell}:\mathcal{S}_{\ell}=\left\{k_{1},\ldots,k_{\ell}\right\}\subseteq\mathcal{J}_{-j}}\frac{\exp\left(-v_{j}\right)}{\sum_{k\in\mathcal{S}_{\ell,j}}\exp\left(-v_{k})\right)}\ln\left(\sum_{k\in\mathcal{S}_{\ell,j}}\exp\left(-v_{k})\right)\right)+\left(\sum_{j=1}^{J}v_{j}-\gamma\right)$
where the third and last equalities use
@@
Proof of Proposition 4. Let $\mathbf{1}_{J}$ be the $J\times 1$ vectors of ones. Since the choice probabilities do not change when a constant $\mu\in\mathbb{R}$ is added to each $V_{ij}\left(\beta\right)$, $E\sum_{k=1}^{J}Y_{ik}\log P^{\mathrm{Q}}(Y_{i}=k|\mathcal{J},V_{i}\left(\beta\right)+\mu\mathbf{1}_{J})$ is a constant function of $\mu.$ This implies that, as a function of $\mu,$ the derivative of $E\sum_{k=1}^{J}Y_{ik}\log P^{\mathrm{Q}}(Y_{i}=k|\mathcal{J},V_{i}\left(\beta\right)+\mu\mathbf{1}_{J})$ with respect to $\mu$ is a zero function for any $V_{i}\left(\beta\right).$ This derivative at $\mu=0$ is
@@
Hence, $E\left[\mathbf{1}_{J}^{\prime}\Lambda_{i}^{\mathrm{Q}}\left(X_{i}\beta\right)\overrightarrow{Y}_{i}\right]=0$ for any $\beta\in\mathcal{B}$. In particular, $E\left[\mathbf{1}_{J}^{\prime}\Lambda_{i}^{\mathrm{Q}}\left(X_{i}\beta_{\mathrm{Q}}^{\ast}\right)\overrightarrow{Y}_{i}\right]=0.$ Combining this with Assumption (i) of the Proposition, we conclude that $\beta=\beta_{\mathrm{Q}}^{\ast}$ is the unique solution to
@@
Since $\beta_{0}\neq 0,$ we know that $\beta_{0,\ell}\neq 0$ for some $\ell=1,\ldots,$ or $L.$ Without loss of generality, we assume that $\beta_{0,1}\neq 0.$ We partition $X_{i}$ into $X_{i}=(X_{i}^{(1)},X_{i}^{(2)})\in\mathbb{R}^{J\times L}$ with $X_{i}^{(1)}\in\mathbb{R}^{J\times 1}$ and $X_{i}^{(2)}\in\mathbb{R}^{J\times\left(L-1\right)}.$ For any $\beta\in\mathbb{R}^{L},$ partition it conformably as $\beta=\left(\beta_{1},\beta_{2}^{\prime}\right)^{\prime}$ where $\beta_{1}\in\mathbb{R}$ and $\beta_{2}\in\mathbb{R}^{L-1}.$ In particular, we write $\beta_{0}=\left(\beta_{0,1}^{\prime},\beta_{0,2}^{\prime}\right)^{\prime}$ and $\beta_{\mathrm{Q}}^{\ast}=\left(\beta_{\mathrm{Q},1}^{\ast\prime},\beta_{\mathrm{Q},2}^{\ast\prime}\right)^{\prime},$ where $\beta_{0,1}$ and $\beta_{\mathrm{Q},1}^{\ast}$ are the first elements of $\beta_{0}$ and $\beta_{\mathrm{Q}}^{\ast},$ respectively. Define $\delta_{1}=\beta_{1}/\beta_{0,1}$ and $\delta_{2}=\beta_{2}-\delta_{1}\beta_{0,2}.$ Then, we have the reparametrization:
@@
where $\mathcal{X}_{i}=(\mathcal{X}_{i}^{\left(1\right)},\mathcal{X}_{i}^{\left(2\right)})$, $\mathcal{X}_{i}^{(1)}=\left(X_{i}\beta_{0}\right)\in\mathbb{R}^{J\times 1},$ and $\mathcal{X}_{i}^{(2)}=X_{i}^{\left(2\right)}\in\mathbb{R}^{J\times\left(L-1\right)}.$ Note that
@@
$=\Theta_{0}E\left\{\boldsymbol{1}_{J}^{\prime}\Lambda_{i}^{\mathrm{Q}}\left(\mathcal{X}_{i}^{(1)}\delta_{1}^{\circ}\right)\overrightarrow{Y}_{i}\right\}+\Theta_{1}E\left\{\left[\mathcal{X}_{i}^{(1)}\right]^{\prime}\Lambda_{i}^{\mathrm{Q}}\left(\mathcal{X}_{i}^{(1)}\delta_{1}^{\circ}\right)\overrightarrow{Y}_{i}\right\}=\mathbf{0}.$
We have therefore shown that $\delta=\left(\delta_{1}^{\circ},\mathbf{0}_{1\times\left(L-1\right)}\right)^{\prime}$ is a solution to (23).
@@
$\delta_{1}^{\ast}=\delta_{1}^{\circ}\text{ and }\delta_{2}^{\ast}=\beta_{\mathrm{Q},2}^{\ast}-\frac{\beta_{\mathrm{Q},1}^{\ast}}{\beta_{0,1}}\beta_{0,2}=\mathbf{0}.$

### v2

@@ 3.4 Patterns of Substitution and the Lack of IIA
We note in passing that IIA also does not hold for standard multinomial probit models, i.e., the multinomial probit with iid normal random utility components. See Section S.D in the supplementary appendix for details.
@@ 4.2 On the Estimation of Parameter Ratios
Assume that (i) $\hat{\beta}_{\mathrm{Q}}\rightarrow^{p}\beta_{\mathrm{Q}}^{\ast}\in Int(\mathcal{B)}$, and $\beta_{\mathrm{Q}}^{\ast}$ is the unique solution to $E\left[X_{i}^{\prime}\Lambda_{i}^{\mathrm{Q}}\left(X_{i}\beta\right)\overrightarrow{Y}_{i}\right]=0;$
@@ 8 Conclusions and Further Research Directions
Third, there are various other flavors of logit models used by economists. The nested logit explicitly allows a correlation between different subsets of alternatives that are uncorrelated within a nest. Exploring the properties of the SEVI analogue would be of interest, given the quite different patterns of substitution within a nest. The generalized multinomial logit’s gamma parameter determines how the random component of the mixed logit interacts with the scale parameter (Fiebig et al., (2010)). It is unclear how this tradeoff is influenced by a SEVI or LEVI core. Another interesting question is how the cut-off points and parameter estimates are influenced in an ordered logit model when switching its foundation from LEVI to SEVI error components. As discussed in Section 3.3, the SEVI analogue of the LEVI-type model for completely rank-ordered can be developed, but the exact details warrant further studies.
@@ 9 Appendix of Proofs
Proof of Theorem 1. A direct algebraic proof is given in the supplementary appendix (see Section S.A). Here, we prove the result using the inclusion-exclusion principle. For each $k\in\mathcal{J}_{-j}:=\left\{1,\ldots,J\right\}\backslash\left\{j\right\}$, define the event
$\mathcal{E}_{k,j}:v_{ij}+\varepsilon_{ij}\geq v_{ik}+\varepsilon_{ik},$
where $\varepsilon_{ij}\thicksim iid$ $\mathrm{SEVI}$ over $j\in\mathcal{J}$. Equivalently, we define the event $\mathcal{E}_{k,j}$ as
$\mathcal{E}_{k,j}:-v_{ij}+\tilde{\varepsilon}_{ij}\leq-v_{ik}+\tilde{\varepsilon}_{ik},$
@@
$=1-\sum_{k_{1}:\left\{k_{1}\right\}\subseteq\mathcal{J}_{-j}}\Pr\left(\mathcal{E}_{k,j}^{c}|v_{i}\right)+\sum_{k_{1}<k_{2}:\left\{k_{1},k_{2}\right\}\subseteq\mathcal{J}_{-j}}\Pr\left(\mathcal{E}_{k_{1},j}^{c}\cap\mathcal{E}_{k_{2},j}^{c}|v_{i}\right)$
$+\ldots+\left(-1\right)^{J-1}\sum_{k_{1}<\ldots<k_{J-1}:\left\{k_{1},\ldots,k_{J-1}\right\}\subseteq\mathcal{J}_{-j}}\Pr\left(\mathcal{E}_{k_{1,j}}^{c}\cap\mathcal{E}_{k_{2},j}^{c}\cap\ldots\cap\mathcal{E}_{k_{J-1},j}^{c}|v_{i}\right),$ (21)
where we have used the inclusion-exclusion principle (e.g., Proposition 4.4 of Ross, (2013)) for computing $\Pr(\cup_{{}_{k\in\mathcal{J}_{-j}}}\mathcal{E}_{k,j}^{c}|v_{i})$, the probability of a union of $\left(J-1\right)$ events.
@@
$-v_{ij}+\tilde{\varepsilon}_{ij}>-v_{ik}+\tilde{\varepsilon}_{ik}\text{ for all }k\in\left\{k_{1},\ldots,k_{\ell}\right\}\subseteq\mathcal{J}_{-j}.$
The probability of the above event is the same as the probability of choosing the $j$-th alternative out of the set of alternatives $\left\{k_{1},\ldots,k_{\ell},j\right\}$ in the standard LEVI model with $\left\{-v_{ik_{1}},\ldots,-v_{ik_{\ell}},-v_{ij}\right\}$ as the systematic utilities. Hence, this probability is equal to
@@
Proof of Proposition 2. The first result is the same as that given in Theorem 1. To prove the second result, we note that
@@
Proof of Proposition 3. Recall that $G\left(\cdot\right)$ and $g\left(\cdot\right)$ are the CDF and the pdf of the LEVI distribution. For $\varepsilon_{j}\thicksim iid$ $\mathrm{SEVI}$ and $\tilde{\varepsilon}_{j}\thicksim iid$ $\mathrm{LEVI},$ we have
@@
$=\sum_{j=1}^{J}E\left[v_{j}+\varepsilon_{j}|v_{j}+\varepsilon_{j}\geq v_{k}+\varepsilon_{k},\text{ }k\in\mathcal{J}_{-j}\right]P^{\mathrm{SEVI}}\left(Y=j|v\right)$
@@
$=\sum_{j=1}^{J}\left[v_{j}-E\left(\tilde{\varepsilon}_{j}|v_{k}-v_{j}+\tilde{\varepsilon}_{j}\leq\tilde{\varepsilon}_{k},\text{ }k\in\mathcal{J}_{-j}\right)\right]P^{\mathrm{SEVI}}\left(Y=j|v\right)$
@@
By employing a change of variable $(s=\exp(-z))$ and letting
@@
$=\left[\gamma+v_{j}+\ln\left(\sum_{k\in\mathcal{S}_{\ell}\cup\left\{j\right\}}\exp\left(-v_{k})\right)\right)\right]\frac{\exp\left(-v_{j}\right)}{\sum_{k\in\mathcal{S}_{\ell}\cup\left\{j\right\}}\exp\left(-v_{k})\right)}.$
where the second-to-last line follows because for any $\xi>0,$
@@
[…]
For notational simplicity, let $\mathcal{S}_{\ell,j}:=\mathcal{S}_{\ell}\cup\left\{j\right\}.$ Then
@@
$=\left(\sum_{j=1}^{J}v_{j}-\gamma\right)-\sum_{j=1}^{J}\sum_{\ell=1}^{J-1}\left(-1\right)^{\ell}\sum_{k_{1}<\ldots<k_{\ell}:\mathcal{S}_{\ell}=\left\{k_{1},\ldots,k_{\ell}\right\}\subseteq\mathcal{J}_{-j}}\frac{\exp\left(-v_{j}\right)}{\sum_{k\in\mathcal{S}_{\ell,j}}\exp\left(-v_{k})\right)}\ln\left(\sum_{k\in\mathcal{S}_{\ell,j}}\exp\left(-v_{k})\right)\right),$
where the third and last equalities use the result:
@@
Proof of Proposition 4. Let $\mathbf{1}_{J}$ be the $J\times 1$ vector of ones. Since the choice probabilities do not change when a constant $\mu\in\mathbb{R}$ is added to each $V_{ij}\left(\beta\right)$, $E\sum_{k=1}^{J}Y_{ik}\log P^{\mathrm{Q}}(Y_{i}=k|\mathcal{J},V_{i}\left(\beta\right)+\mu\mathbf{1}_{J})$ is a constant function of $\mu.$ This implies that, as a function of $\mu,$ the derivative of $E\sum_{k=1}^{J}Y_{ik}\log P^{\mathrm{Q}}(Y_{i}=k|\mathcal{J},V_{i}\left(\beta\right)+\mu\mathbf{1}_{J})$ with respect to $\mu$ is a zero function for any $V_{i}\left(\beta\right).$ But this derivative at $\mu=0$ is
@@
Hence, $E\left[\mathbf{1}_{J}^{\prime}\Lambda_{i}^{\mathrm{Q}}\left(X_{i}\beta\right)\overrightarrow{Y}_{i}\right]=0$ for any $\beta\in int(\mathcal{B)}$. In particular, $E\left[\mathbf{1}_{J}^{\prime}\Lambda_{i}^{\mathrm{Q}}\left(X_{i}\beta_{\mathrm{Q}}^{\ast}\right)\overrightarrow{Y}_{i}\right]=0.$ Combining this with Assumption (i) of the Proposition, we conclude that $\beta=\beta_{\mathrm{Q}}^{\ast}$ is the unique solution to
@@
Since $\beta_{0}\neq 0,$ we know that $\beta_{0,\ell}\neq 0$ for some $\ell=1,\ldots,$ or $L.$ Without loss of generality, we assume that $\beta_{0,1}\neq 0.$ We partition $X_{i}$ into $X_{i}=(X_{i}^{(1)},X_{i}^{(2)})\in\mathbb{R}^{J\times L}$ with $X_{i}^{(1)}\in\mathbb{R}^{J\times 1}$ and $X_{i}^{(2)}\in\mathbb{R}^{J\times\left(L-1\right)}.$ For any $\beta\in\mathbb{R}^{L},$ partition it conformably as $\beta=\left(\beta_{1},\beta_{2}^{\prime}\right)^{\prime}$ where $\beta_{1}\in\mathbb{R}$ and $\beta_{2}\in\mathbb{R}^{L-1}.$ In particular, we write $\beta_{0}=\left(\beta_{0,1}^{\prime},\beta_{0,2}^{\prime}\right)^{\prime}$ and $\beta_{\mathrm{Q}}^{\ast}=\left(\beta_{\mathrm{Q},1}^{\ast\prime},\beta_{\mathrm{Q},2}^{\ast\prime}\right)^{\prime},$ where $\beta_{0,1}$ and $\beta_{\mathrm{Q},1}^{\ast}$ are the first elements of $\beta_{0}$ and $\beta_{\mathrm{Q}}^{\ast},$ respectively. Define $\delta_{1}=\beta_{1}/\beta_{0,1}$ and $\delta_{2}=\beta_{2}-\delta_{1}\beta_{0,2}.$ Then, we have the following reparametrization:
@@
$=\left[X_{i}\beta_{0}-X_{i}^{\left(1\right)}\beta_{0,1}-X_{i}^{\left(2\right)}\beta_{0,2}\right]+\left[X_{i}^{\left(1\right)}\beta_{0,1}+X_{i}^{\left(2\right)}\beta_{0,2}\right]\delta_{1}+X_{i}^{\left(2\right)}\delta_{2}$
@@
where $\mathcal{X}_{i}=(\mathcal{X}_{i}^{\left(1\right)},\mathcal{X}_{i}^{\left(2\right)})$, $\mathcal{X}_{i}^{(1)}=X_{i}\beta_{0}\in\mathbb{R}^{J\times 1},$ and $\mathcal{X}_{i}^{(2)}=X_{i}^{\left(2\right)}\in\mathbb{R}^{J\times\left(L-1\right)}.$ Note that
@@
$E\left\{\left[\mathcal{X}_{i}^{(2)}\right]^{\prime}\Lambda_{i}^{\mathrm{Q}}\left(\mathcal{X}_{i}^{\left(1\right)}\delta_{1}+\mathcal{X}_{i}^{\left(2\right)}\delta_{2}\right)\overrightarrow{Y}_{i}\right\}_{\delta_{1}=\delta_{1}^{\circ},\text{ }\delta_{2}=\mathbf{0}}$
@@
$=\Theta_{0}E\left\{\boldsymbol{1}_{J}^{\prime}\Lambda_{i}^{\mathrm{Q}}\left(\mathcal{X}_{i}^{(1)}\delta_{1}^{\circ}\right)\overrightarrow{Y}_{i}\right\}+\Theta_{1}E\left\{\left[\mathcal{X}_{i}^{(1)}\right]^{\prime}\Lambda_{i}^{\mathrm{Q}}\left(\mathcal{X}_{i}^{(1)}\delta_{1}^{\circ}\right)\overrightarrow{Y}_{i}\right\}=0.$
We have therefore shown that $\delta=\left(\delta_{1}^{\circ},0_{1\times\left(L-1\right)}\right)^{\prime}$ is a solution to (23).
@@
$\delta_{1}^{\ast}=\delta_{1}^{\circ}\text{ and }\delta_{2}^{\ast}=\beta_{\mathrm{Q},2}^{\ast}-\frac{\beta_{\mathrm{Q},1}^{\ast}}{\beta_{0,1}}\beta_{0,2}=0.$